In the following table, Pt represents the price at time t and Qt represents shares outstanding at time t. If the equally weighted index of the three stocks at t=0 was 300, what is the index at t=1? Please show calculations.

P0 Q0 P1 Q1
A 88 100 93 100
B 45 200 40 200
C 96 200 48 400
Answer #1

For calculating an index, we use following formula

Index Value = \sum (Market Cap of securities in index) / Index Divisor

Index Divisor remains constant. It changes in case securities are added/dropped from the index

AT t = 0

А в с Price Shares Market Cap = Price * Shares I 881 100=B3* C3 no

SUMIE X v fx =D3/$D$6 AA B C D Price Weighting = Market Cap of Shares Market Cap = Price * Shares security / Total Market Cap

Price Shares Market Cap = Price * Shares Weighting = Market Cap of security / Total Market Cap
88 100 8800 24%
45 200 9000 24%
96 200 19200 52%
Total 500 37000 100%

Hence,

Index Value = \sum (Market Cap of securities in index) / Index Divisor

300 = 37000 / Index Divisor

Index Divisor = 37000 / 300 = 123.333333

Now,

Using this for T = 1

AA в с Price Weighting = Market Cap of Shares Market Cap = Price * Shares security / Total Market Cap 100 9300| 25% 2001 8000

E4 ® =D4/$D$6 fx C A B D E F Price Weighting = Market Cap of Shares Market Cap = Price * Shares security / Total Market Cap 1

Hence,

Index Value = \sum (Market Cap of securities in index) / Index Divisor

Index Value = 36500 / 123.33333

Index Value = 295.9459459 = 295.95